a2+b2+c2=-(ab+bc+ca) find the them 3a-2b+5c value of what ?

Grade:12th pass

3 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please attach the image of the question

Regards
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please upload an image.
We will happy to help you.
Good luck 
Cheers 
Vishal Dagur
one year ago
Given A2+b2+C2= -(ab+bc+ca) then find 3a-2b+5c 
 
So, multiple both side 2
2a2 +2b2 +2c= -2ab-2bc-2ca
 
We right -- 2a= a2 +a,2b2=b2 +b,2c2=c2+c2
So, 
 
a2+a2+b2+b2+c2+c2+2ab+2bc+2ca=0
 
This make formula 
(A+B)2+(B+C)2+(C+A)2=0
 
(A+B)2=0 ,A+B =0 ,so A= -B
(B+C)2=0 , B=-C
(C+A)2=0 , C= -A
 
Now 
3a-2b+5c. 
Put a =-B  and c=-A
-3b-2b+5(-A).  (We know a=-B so  -A =b)
 -3b-2b+5b
-5b+5b = 0. 
 
So there answer is zero